OpenStudy (anonymous):

Let f(x) = 5x – 26 and g(x) = 6x. Part 1 [4 points] What is (f ◦ g)(x) ? Part 2 [2 points] Use complete sentences to describe the method you used to solve this problem.

OpenStudy (anonymous):

(f ◦ g)(x) means f(g(x)) f(x)=5x-26 f(g(x))=5g(x)-26=5*6x-26=30x-26

OpenStudy (anonymous):

(f*g)(x) 5(6x)-26 30x-26 Use the subsitution method, then expand it.

OpenStudy (anonymous):

does that answer it completly ?